Quiz 2 part 1 (10 points). Using the following income statement and balance sheet prepare the financial statement of cash flow (format it on the next page) and submit it via dropbex by 01/28/2018 Nabors, Inc. 2011 Income Statement (S in millions) $9,610 6,310 Net sales Less: Cost of goods sold Less: Depreciation Earnings before interest and taxes Less: Interest paid Taxable Income Less: Taxes Net income 370 1,930 $1,300 $ 845 Nabors, Inc 2010 and 2011 Balance Sheets (S in millions) 2010 2011 2011 S 310 S 405 2,640 3,055 2010 Accounts payable Notes payable S 2,720 S 2,570 Cash Accounts ree. Inventory Total Net fixed assets 3,2753.850 Total S6225 S 7310Long-term debt 7,875 5,000 」-490 8,100 5,250 」,060 S17.980 」0960-10670 Common stock Retained earnings Total asset S17,185 S17,980 Total liab.& equity S17185 Note: Retained carnings.is listed as a separate line item in this balance sheet (where as retained earnings were part of common stock in the balance sheet before)Use the information provided to answer the questions regarding Nabors, Inc.
Answer #1
1) STATEMENT OF CASH FLOWS
Cash flows from operating activities:
Net income 845
Adjustments to reconcile net income with net cash from operating activities:
Depreciation 1370
Increase in accounts receivable -415
Increase in inventory -575
Decrease in accounts payable -150 230
Net cash provided by operating activities 1075
Cash flows from investing activities:
Purchase of equipment (10670-10960+1370) -1080
Net cash used by operating activities -1080
Cash flows from financing activities:
Increase in long term debt (8100-7875) 225
Redemption of notes payable -100
Dividends paid (2060-1490-845) -275
Common stock issued 250
Net cash used by financing activitie 100
Net increase in cash and cash equivalents 95
Cash & cash equivalents at the beginning of year 310
Cash & cash equivalents at the end of year 405
2)a) CASH FLOW FROM ASSETS (Using the formulae given)
Operating cash flow = 1930+1370-455= 2845
Net capital spending = 10670-10960+1370 = 1080
Changes in NWC = (7310-2570)-(6225-2720) = 1235
CASH FLOW FROM ASSETS (2845-1080-1235) = 530
2)b) CASH FLOW TO CREDITORS = 630-125 = 505
2)c) CASH FLOW TO STOCKHOLDERS = 275-250 = 25